A body weighs 250 16, at the earths surface. Determine: a) The mass of the body. b) The weight of the body on the moons cu
Answer #1

weight of the body

1 lb = 4.4482216282509 N

w = 250 lb= 1112.05540706273 N

m = w/g = 1112.05540706273 N/9.8 m/s^2 = 113.47 kg or 7.77 slug

(b)

weight of the body on moon

w= mg _ moon = 113.47 kg (9.8/6) = 185.34 N or 41.66 lb